# Matchings in hypercubes extend to long cycles

### v1

@@ 2.4. Large cuts through maximal matchings
Suppose that $M$ contains a long edge $uv$, i.e., $\ell(uv)\geq 2$. We define $M^{\prime}:=M\setminus\{uv\}$. Furthermore, if there is a neighbor $u^{\prime}$ of $u$ in $Q_{d}$ that is not covered by $M^{\prime}$, then we define $M^{\prime}:=M^{\prime}\cup\{uu^{\prime}\}$. Similarly, if there is a neighbor $v^{\prime}$ of $v$ in $Q_{d}$ that is not covered by $M^{\prime}$, then we define $M^{\prime}:=M^{\prime}\cup\{vv^{\prime}\}$. As an edge of length $\ell(uv)\geq 2$ is removed from $M$ and at most two length 1 edges are added instead, we have $\ell(M)\geq\ell(M^{\prime})$. Furthermore, note that $M^{\prime}$ is again a maximal matching with one less long edge than $M$. Consequently, repeating this replacement process terminates with a maximal matching of $Q_{d}$. This completes the proof of the lemma. ∎
@@
For dimension $d=6$, Havel and Křivánek [HK82] improved the lower bound of $\lceil f(6)\rceil=23$ guaranteed by Lemma 19 to $24$, but we do not need this strengthening for our arguments.
@@
Let $M$ be matching of $K(Q_{d})$ that avoids $z=\emptyset$ and satisfies property (H) in Theorem 8. Let $i\in[d]$ and $u,u^{i}$ with $u\in V(Q^{i}_{0})\setminus\{z\}$ be two vertices avoided by $M$. Then $M\cup\{uu^{i}\}$ violates property (H) if and only if one of the following two conditions holds:
@@
Statements (i) and (ii) follow directly from the definition of property (H). The next claim follows from the fact that $V(Q^{i}_{0})\setminus\{z,u\}$ has even cardinality, and every edge in $M^{i}_{-}$ covers exactly one of the vertices from this set, whereas every edge in $M^{i}_{0}$ covers two of them. To prove the last part we argue as follows: If (i) holds, then $M$ contains $2^{d-2}$ edges in the half-layer in direction $i$, which all belong to $M^{i}_{-}$, plus at least $(2^{d-2}-2)/2=2^{d-3}-1$ edges that cover all remaining vertices in $Q^{i}_{0}$ except $z$ and $u$, which in total gives $13\cdot 2^{d-3}-$ 1 edges in $M$. If (ii) holds, then $M$ contains $12^{d-2}-$ 1 edges in the near half-layer in direction $i$, which all belong to $M^{i}_{-}$, plus at least $\lceil(2^{d-2}-1)/2\rceil=2^{d-3}$ edges that cover all remaining vertices in $Q^{i}_{0}$ except $z$ and $u$, at least one of which belongs to $M^{i}_{-}$, which in total gives $2^{d-2}$ edges in $M^{i}_{-}$ and $13\cdot 2^{d-3}-$ 1 edges in $M$. ∎
@@
The first case is that there are vertices $u,u^{i}$ with $i\in[d]$ that are avoided by $M$ and different from $z$. In this case $M\cup\{uu^{i}\}$ violates property (H), and applying Lemma 22 yields $|M^{i}_{-}|\geq 2^{d-2}$. For $d=5$ and $d\geq 6$ this lower bound evaluates to 8 and $\geq 16$, respectively.
@@
The third case is that one of the neighbors $u$ of $z$ in $Q_{d}$ is avoided by $M$ and $M\cup\{uz\}$ is maximal. In this case we obtain $\ell(M)\geq f(d)-1$, and consequently there is a direction $i\in[d]$ with $|M^{i}_{-}|\geq\lceil(f(d)-1)/d\rceil$. For $d=5$ and $d\geq 6$ the function $\lceil(f(d)-1)/d\rceil$ evaluates to 3 and $\geq 4$, respectively.
@@ 3. Proof of Theorem 5
For $2\leq d\leq 4$, we verified the theorem with computer help.
@@
If $M$ is perfect, we apply Theorem 2. Otherwise, if $M$ has no half-layers in any direction, then we choose a vertex $u$ avoided by $M$. Otherwise, if $M$ contains a half-layer in a direction $i\in[d]$ such that $M$ avoids at least two vertices of $Q^{i}_{0}$ or $Q^{i}_{1}$, then we choose a vertex $u$ avoided by $M$ in $Q^{i}_{0}$ or $Q^{i}_{1}$, respectively. By Lemma 15 (i) and the fact that $M$ is not perfect, $M$ contains no other half-layer. Consequently, in the latter two cases, $M$ and $u$ satisfy property (H). We may thus apply Theorem 8 to obtain a cycle that extends $M$ (and avoids $u$).
In the remaining case, $M$ contains a half-layer in a direction $i\in[d]$ and avoids exactly one vertex $u$ in $Q^{i}_{0}$ and one vertex $v$ in $Q^{i}_{1}$. If $v=u^{i}$, we apply Theorem 2 to the perfect matching $M\cup\{uv\}$. Otherwise consider the modified matching $N:=M\setminus\{u^{i}u^{iM}\}\cup\{uu^{iM}\}$, which avoids two vertices in $Q^{i}_{1}$, namely $u^{i}$ and $v$. Consequently, by Theorem 8 there is a cycle $C$ that extends $N$ and avoids $u^{i}$. Then, the cycle $C\setminus\{uu^{iM}\}\cup\{uu^{i},u^{i}u^{iM}\}$ extends $M$. ∎
@@ 4.1. Forward implication
Proof of Theorem 8 (normal-$\Rightarrow$).
For the sake of contradiction suppose that $M$ violates property (H) and that there is a cycle $C$ that extends $M$ and avoids $z$. Then for some $i\in[d]$ there is a half-layer in direction $i$ and all vertices of $Q^{i}_{0}$ except $z$ are covered by $M$. Every odd vertex of $Q^{i}_{0}$ is end vertex of an edge of the half-layer, and is therefore connected in $C\setminus M$ to an even vertex of $Q^{i}_{0}$. As every vertex except $z$ is covered by $M$, no two of these even vertices are the same. It follows that $C$ visits all odd vertices of $Q^{i}_{0}$ and at least as many even vertices, but at the same time it avoids the even vertex $z$, a contradiction. ∎
@@ 4.2. Reverse implication (induction step $d-1\rightarrow d$ for $d\geq 6$)
Proof of Theorem 8 (normal-$\Leftarrow$) for $d\geq 6$.
@@
Case 1: $|M^{i}_{-}|\geq 4$ is even. By (i), we can apply Lemma 17 and obtain a perfect matching $P_{0}$ on $K(A_{0})$ such that $M^{i}_{0}\cup P_{0}$ contains no half-layers. Applying Theorem 8 inductively to $K(Q^{i}_{0})$, we obtain that $M^{i}_{0}\cup P_{0}$ can be extended to a cycle $C_{0}$ that avoids $z$. Let $P_{1}$ be the shortcut edges of the linear forest $(C_{0}\setminus P_{0})\cup M^{i}_{-}$, which all belong to $K(Q^{i}_{1})$ and form a perfect matching on $K(A_{1})$, where $A_{1}:=V(Q^{i}_{1})\cap V(M^{i}_{-})$. Applying Theorem 5 inductively to $K(Q^{i}_{1})$, we obtain that $M^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$. Observe that $C:=(C_{0}\setminus P_{0})\cup M^{i}_{-}\cup(C_{1}\setminus P_{1})$ is a single cycle in $K(Q_{d})$ that extends $M$ and avoids $z$, as desired.
Case 2: $|M^{i}_{-}|\geq 5$ is odd. We first give an outline of the construction steps in this case. Several details are filled in subsequently, and we also later verify that all assumptions needed to apply the various theorems are indeed satisfied. The points with missing information in the outline are labeled 14.
@@
The construction of $C$ proceeds as follows: We carefully choose a perfect matching $P_{0}$ on $K(B_{0})$, where $B_{0}:=V(Q^{i}_{0})\cap V(N^{i}_{-})$ 2. Having chosen $u$ and $P_{0}$, we apply Theorem 8 inductively to $K(Q^{i}_{0})$ and obtain that $N^{i}_{0}\cup P_{0}$ can be extended to a cycle $C_{0}$ that avoids $z$. If $u\in V(M)$, then we argue that $C_{0}$ also avoids $u$ 3. Let $P_{1}$ be the shortcut edges of the linear forest $(C_{0}\setminus P_{0})\cup N^{i}_{-}$. If $u^{i}\notin V(M)$, then we apply Theorem 5 and obtain that $N^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$. Otherwise, we apply Theorem 8 and obtain that $N^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$ that avoids $u^{i}$. As mentioned before, we need to argue that the assumptions of the those theorems are met in the subcubes $Q^{i}_{0}$ and $Q^{i}_{1}$ for the matchings $N^{i}_{0}\cup P_{0}$ and $N^{i}_{1}\cup P_{1}$, and this is where our choices of $u$ and $P_{0}$ become crucial 4. Observe that $C:=(C_{0}\setminus P_{0})\cup N^{i}_{-}\cup(C_{1}\setminus P_{1})$ is a single cycle that extends $N$ and avoids $z$ and $u$ if $u\in V(M)$ as well as $u^{i}$ if $u^{i}\in V(M)$.
After giving this outline, we now provide the missing details for the points 14.
@@
Case aii: If $X_{01}=\emptyset$, then we have $X_{0}=X_{00}$. As this set has even cardinality and is nonempty, we have $|X_{00}|\geq 2$, i.e., there are at least two distinct vertices $x,y\in X_{0}$ with $x^{iM},y^{iM}\in V(Q^{i}_{0})$. From (i) we know that $M^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We argue that (at least) one of $M^{i}_{0}\cup\{xx^{iM}\}$ or $M^{i}_{0}\cup\{yy^{iM}\}$ contains no (covered near) half-layers of $Q^{i}_{0}$, which implies that we can define $u:=x$ or $u:=y$, respectively. Suppose for the sake of contradiction that this is false, i.e., both $M^{i}_{0}\cup\{xx^{iM}\}$ and $M^{i}_{0}\cup\{yy^{iM}\}$ contain a (covered near) half-layer of $Q^{i}_{0}$. As $M^{i}_{0}\cup\{xx^{iM}\}$ contains a (covered near) half-layer, we obtain that $M^{i}_{0}$ contains a 2-near half-layer $L_{x}$ of $Q^{i}_{0}$, where $x$ is one of the extension vertices and the remaining three extension vertices are covered by $M$; see Figure 5. Symmetrically, as $M^{i}_{0}\cup\{yy^{iM}\}$ contains a (covered near) half-layer, we obtain that $M^{i}_{0}$ contains a 2-near half-layer $L_{y}$ of $Q^{i}_{0}$, where $y$ is one of the extension vertices and the remaining three extension vertices are covered by $M$. As all extension vertices of $L_{x}$ and $L_{y}$ except $x$ or $y$, respectively, are covered by $M$, we obtain that $L_{x}\cap L_{y}=\emptyset$. On the other hand, by Lemma 15 (iii), $L_{x}$ and $L_{y}$ have the same direction. This implies that there is a direction $j\in[d]\setminus\{i\}$ such that $|M^{j}_{-}|\geq|L_{x}|+|L_{y}|=2\cdot(2^{d-3}-2)\geq 12$ and $|M^{i}_{-}|\leq 2\cdot 4=8$, and from the end vertices of $M^{i}_{-}$ in $Q^{i}_{0}$, at most 4 have the same parity, i.e., $M^{i}_{-}$ contains no (covered near) quad-layers, which contradicts our rules (1)–(3) for choosing the direction $i$.
@@
4 We have argued that in both cases, $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$.
4 We now argue that $N^{i}_{1}$ and $u^{i}$ satisfy property (H) in Theorem 8 for $Q^{i}_{1}$. Note that $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ (in case ai) or $N^{i}_{1}=M^{i}_{1}$ (in case aii). It follows that if $M^{i}_{1}$ contains no half-layers of $Q^{i}_{1}$, then $N^{i}_{1}$ contains no half-layers either, so property (H) is trivially satisfied. It remains to consider the case that $M^{i}_{1}$ does contain a half-layer $L$ of $Q^{i}_{1}$. Let $j\in[d]\setminus\{i\}$ be the direction of $L$. By (ii), this means that $M^{i}_{-}$ contains a (covered near) quad-layer of $Q_{d}$; see Figure 6. Lemma 15 (i) yields that $M^{i}_{1}$ contains no half-layers of $Q^{i}_{1}$ in a direction different from $j$, and thus $L$ is the only half-layer of $Q^{i}_{1}$ in $M^{i}_{1}$. If $u^{i}\in V(L)$, then $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ contains no half-layers, so property (H) is trivially satisfied. If $u^{i}\notin V(L)$, then the vertex $u^{i}$ together with all vertices in $X_{1}$ lie in either $(Q^{i}_{1})^{j}_{0}$ or $(Q^{i}_{1})^{j}_{1}$. As $|X_{1}|$ is nonempty, $N^{i}_{1}$ and $u^{i}\notin X_{1}$ satisfy property (H).
@@
Case b: $X=\emptyset$. We define $Y:=\{x\in V(Q^{i}_{0})\mid x\text{ is odd and }xx^{i}\notin M\}$, i.e., these are odd vertices in $V(Q^{i}_{0})$ for which the incident edge in direction $i$ is not in $M$. Clearly, we have $z\notin Y$ as $z=\emptyset$ is even. Furthermore, as $X=\emptyset$ and $M$ is assumed to satisfy property (H), we know that $Y\neq\emptyset$.
@@
    If $M^{i}_{1}$ contains a near half-layer $L^{\prime}$ of $Q^{i}_{1}$, then we select $u\in Y$ as a vertex such that $u^{i}$ is an (even) end vertex of one of edges of $L^{\prime}$;
@@
4 We now argue that $N^{i}_{0}$ contains no (covered near) half-layer of $Q^{i}_{0}$. Note that if $N^{i}_{0}=M^{i}_{0}$ or $N^{i}_{0}=M^{i}_{0}\setminus\{uu^{M}\}$, then the claim follows directly from (i). It remains to consider the case $N^{i}_{0}=M^{i}_{0}\setminus\{uu^{M}\}\cup\{u^{M}u^{iM}\}$, which occurs if and only if $u^{M},u^{iM}\in V(Q^{i}_{0})$. In particular, rule (1’) does not apply, as it would imply $u^{iM}\in V(Q^{i}_{1})$. If $N^{i}_{0}$ contains a (covered near) half-layer $L^{\prime}$, then by (i) $u^{M}u^{iM}$ is one of its edges. Therefore $M^{i}_{0}$ contains the 2-near half-layer $L^{\prime\prime}:=L^{\prime}\setminus\{u^{M}u^{iM}\}$; see Figure 7 (a). Let $j\in[d]\setminus\{i\}$ be the direction of $L^{\prime\prime}$, which is also the direction of the edge $u^{M}u^{iM}$. By Lemma 15 (iii), $M^{i}_{0}$ contains no 2-near half-layers of $Q^{i}_{0}$ in a direction different from $j$, so only $L^{\prime\prime}$ and possibly a second 2-near half-layer in direction $j$ qualify for application of rule (2’). However, the edge $uu^{M}$ removed by rule (2’) also has the direction $j$. This is a contradiction, however, as the edges $uu^{M}$ and $u^{M}u^{iM}$ have different directions (as $u\neq u^{iM}$).

### v2

@@ 2.4. Large cuts through maximal matchings
Suppose that $M$ contains a long edge $uv$, i.e., $\ell(uv)\geq 2$. We define $M^{\prime}:=M\setminus\{uv\}$. Furthermore, if there is a neighbor $u^{\prime}$ of $u$ in $Q_{d}$ that is not covered by $M^{\prime}$, then we redefine $M^{\prime}:=M^{\prime}\cup\{uu^{\prime}\}$. Similarly, if there is a neighbor $v^{\prime}$ of $v$ in $Q_{d}$ that is not covered by $M^{\prime}$, then we redefine $M^{\prime}:=M^{\prime}\cup\{vv^{\prime}\}$. As an edge of length $\ell(uv)\geq 2$ is removed from $M$ and at most two length 1 edges are added instead, we have $\ell(M)\geq\ell(M^{\prime})$. Furthermore, note that $M^{\prime}$ is again a maximal matching of $K(Q_{d})$ with one less long edge than $M$. Consequently, repeating this replacement process terminates with a maximal matching of $Q_{d}$. This completes the proof of the lemma. ∎
The next two lemmas establish lower bounds for the size of maximal matchings of $Q_{d}$ and $K(Q_{d})$, respectively.
@@
The function $f(d)$ defined in Lemma 19 will be used in several computations in the rest of this section. For dimension $d=6$, Havel and Křivánek [HK82] improved the lower bound of $\lceil f(6)\rceil=23$ guaranteed by Lemma 19 to $24$, but we do not need this strengthening for our arguments.
@@
The next lemma describes how adding an edge to a matching of $K(Q_{d})$ can violate property (H).
@@
Let $d\geq 5$, and let $M$ be matching of $K(Q_{d})$ that avoids $z=\emptyset$ and satisfies property (H) in Theorem 8. Furthermore, let $i\in[d]$ and $u,u^{i}$ with $u\in V(Q^{i}_{0})\setminus\{z\}$ be two vertices avoided by $M$. Then $M\cup\{uu^{i}\}$ violates property (H) if and only if one of the following two conditions holds:
@@
Statements (i) and (ii) follow directly from the definition of property (H). The next claim follows from the fact that $V(Q^{i}_{0})\setminus\{z,u\}$ has even cardinality, and every edge in $M^{i}_{-}$ covers exactly one of the vertices from this set, whereas every edge in $M^{i}_{0}$ covers two of them. To prove the last part we argue as follows: If (i) holds, then $M$ contains $2^{d-2}$ edges in the half-layer in direction $i$, which all belong to $M^{i}_{-}$, plus at least $(2^{d-2}-2)/2=2^{d-3}-1$ edges that cover all remaining vertices in $Q^{i}_{0}$ except $z$ and $u$, which in total gives at least $13\cdot 2^{d-3}-$ 1 edges in $M$. If (ii) holds, then $M$ contains $12^{d-2}-$ 1 edges in the near half-layer in direction $i$, which all belong to $M^{i}_{-}$, plus at least $\lceil(2^{d-2}-1)/2\rceil=2^{d-3}$ edges that cover all remaining vertices in $Q^{i}_{0}$ except $z$ and $u$, at least one of which belongs to $M^{i}_{-}$, which in total gives at least $2^{d-2}$ edges in $M^{i}_{-}$ and at least $13\cdot 2^{d-3}-$ 1 edges in $M$. ∎
@@
The next lemma guarantees us a direction such that many edges of an (H)-maximal matching belong to the layer in this direction. In the proof of Theorem 8 we will choose this direction for splitting the cube into two subcubes and applying induction.
@@
The first case is that there are vertices $u,u^{i}$ with $i\in[d]$ that are avoided by $M$ and different from $z$. In this case $M\cup\{uu^{i}\}$ violates property (H), and applying Lemma 22 yields $|M^{i}_{-}|\geq 2^{d-2}$. For $d=5$ and $d\geq 6$ this lower bound evaluates to 8 and at least 16, respectively.
@@
The third case is that one of the neighbors $u$ of $z$ in $Q_{d}$ is avoided by $M$ and $M\cup\{uz\}$ is maximal. In this case, by Lemmas 18 and 19, we obtain $\ell(M)\geq f(d)-1$, and consequently there is a direction $i\in[d]$ with $|M^{i}_{-}|\geq\lceil(f(d)-1)/d\rceil$. For $d=5$ and $d\geq 6$ the function $\lceil(f(d)-1)/d\rceil$ evaluates to 3 and at least 4, respectively.
@@ 3. Proof of Theorem 5
For $2\leq d\leq 4$, we verified the theorem with computer help. For details, see Section 4.3.
@@
If $M$ is perfect, we apply Theorem 2. Otherwise, if $M$ has no half-layers in any direction, then we choose a vertex $u$ avoided by $M$. Otherwise, if $M$ contains a half-layer in a direction $i\in[d]$ such that $M$ avoids at least two vertices of $Q^{i}_{0}$ or $Q^{i}_{1}$, then we choose a vertex $u$ avoided by $M$ in $Q^{i}_{0}$ or $Q^{i}_{1}$, respectively. By Lemma 15 (i) and the fact that $M$ is not perfect, $M$ contains no other half-layer. Consequently, in the latter two cases, $M$ and $u$ satisfy property (H) (where $u$ plays the role of $z=\emptyset$). We may thus apply Theorem 8 to obtain a cycle that extends $M$ and avoids $u$.
In the remaining case, $M$ contains a half-layer $L$ in a direction $i\in[d]$ and avoids exactly one vertex $u$ in $Q^{i}_{0}$ and one vertex $v$ in $Q^{i}_{1}$. If $v=u^{i}$, we apply Theorem 2 to the perfect matching $M\cup\{uv\}$. Otherwise consider the modified matching $N:=(M\setminus\{u^{i}u^{iM}\})\cup\{uu^{iM}\}$, which avoids two vertices in $Q^{i}_{1}$, namely $u^{i}$ and $v$. By Lemma 15 (i) and the fact that $N$ is not perfect, $N$ contains no half-layer other than $L$. Consequently, $N$ and $u^{i}$ satisfy property (H) (where $u^{i}$ plays the role of $z=\emptyset$), so applying Theorem 8 gives a cycle $C$ that extends $N$ and avoids $u^{i}$. Then, the cycle $(C\setminus\{uu^{iM}\})\cup\{uu^{i},u^{i}u^{iM}\}$ extends $M$. ∎
@@ 4. Proof of Theorem 8
The statement in Theorem 8 is an equivalence, and we prove both directions of the implication in the following subsections. We first consider the forward implication, then the reverse implication for $d\geq 6$, and then the reverse implication for $d=5$ (settled with computer help), followed by some remarks about implementation details for the computer verification.
@@ 4.1. Forward implication
Proof of Theorem 8 ($\Rightarrow$).
For the sake of contradiction suppose that $M$ violates property (H) and that there is a cycle $C$ that extends $M$ and avoids $z$. Then for some $i\in[d]$ there is a half-layer in direction $i$ and all vertices of $Q^{i}_{0}$ except $z$ are covered by $M$. Every odd vertex of $Q^{i}_{0}$ is an end vertex of an edge of the half-layer, and is therefore connected in $C\setminus M$ to an even vertex of $Q^{i}_{0}$. As every vertex of $Q^{i}_{0}$ except $z$ is covered by $M$, no two of these even vertices are the same. It follows that $C$ visits all odd vertices of $Q^{i}_{0}$ and at least as many even vertices, but at the same time it avoids the even vertex $z$, a contradiction. ∎
@@ 4.2. Reverse implication (induction step $d-1\rightarrow d$ for $d\geq 6$)
Proof of Theorem 8 ($\Leftarrow$) for $d\geq 6$.
@@
Case 1: $|M^{i}_{-}|\geq 4$ is even. By (i), we can apply Lemma 17 and obtain a perfect matching $P_{0}$ on $K(A_{0})$ such that $M^{i}_{0}\cup P_{0}$ contains no half-layers. Applying Theorem 8 inductively to $K(Q^{i}_{0})$, we obtain that $M^{i}_{0}\cup P_{0}$ can be extended to a cycle $C_{0}$ that avoids $z$. Let $P_{1}$ be the shortcut edges of the linear forest $(C_{0}\setminus P_{0})\cup M^{i}_{-}$, which all belong to $K(Q^{i}_{1})$ and form a perfect matching on $K(A_{1})$, where $A_{1}:=V(Q^{i}_{1})\cap V(M^{i}_{-})$. Applying Theorem 5 inductively to $K(Q^{i}_{1})$, we obtain that $M^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$. Observe that $C:=(C_{0}\setminus P_{0})\cup M^{i}_{-}\cup(C_{1}\setminus P_{1})$ is a single cycle in $K(Q_{d})$ that extends $M$ and avoids $z$, as desired. Indeed, $C$ is obtained from $C_{1}$ by replacing each edge of $P_{1}$ by the path between the same end vertices starting and ending with an edge from $M^{i}_{-}$ plus edges of $C_{0}\setminus P_{0}$ in between.
Case 2: $|M^{i}_{-}|\geq 5$ is odd. We first give an outline of the construction steps in this case. Several details are filled in subsequently, and we also later verify that all assumptions needed to apply the various theorems are indeed satisfied. The points with missing information in the outline are labeled 15.
@@
The construction of $C$ proceeds as follows: We carefully choose a perfect matching $P_{0}$ on $K(B_{0})$, where $B_{0}:=V(Q^{i}_{0})\cap V(N^{i}_{-})$ 2. Having chosen $u$ and $P_{0}$, we apply Theorem 8 inductively to $K(Q^{i}_{0})$ and obtain that $N^{i}_{0}\cup P_{0}$ can be extended to a cycle $C_{0}$ that avoids $z$. If $u\in V(M)$, then we argue that $C_{0}$ also avoids $u$ 3. Let $P_{1}$ be the shortcut edges of the linear forest $(C_{0}\setminus P_{0})\cup N^{i}_{-}$. If $u^{i}\notin V(M)$, then we apply Theorem 5 and obtain that $N^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$. Otherwise, we apply Theorem 8 and obtain that $N^{i}_{1}\cup P_{1}$ can be extended to a cycle $C_{1}$ that avoids $u^{i}$. As mentioned before, we need to argue that the assumptions of those theorems are met in the subcubes $Q^{i}_{0}$ and $Q^{i}_{1}$ for the matchings $N^{i}_{0}\cup P_{0}$ and $N^{i}_{1}\cup P_{1}$, and this is where our choices of $u$ and $P_{0}$ become crucial; these verifications are done below under the labels 4 and 5 separately for the 0- and 1-subcube, respectively. Observe that $C:=(C_{0}\setminus P_{0})\cup N^{i}_{-}\cup(C_{1}\setminus P_{1})$ is a single cycle that extends $N$ and avoids $z$ and $u$ if $u\in V(M)$ as well as $u^{i}$ if $u^{i}\in V(M)$.
After giving this outline, we now provide the missing details for the points 15. Each of the two steps 1 (choosing $u$) and 2 (choosing $P_{0}$) splits into two cases. The argument for 3 ($C_{0}$ avoids $u$) is straightforward and is presented after step 1, as it requires the definition of $u$. The verifications 4 and 5 are done after step 1 and again after step 2 via some auxiliary claims (in each of the two respective cases), as these two steps are sequential and depend on each other.
@@
Case aii: If $X_{01}=\emptyset$, then we have $X_{0}=X_{00}$. As this set has even cardinality and is nonempty, we have $|X_{00}|\geq 2$, i.e., there are at least two distinct vertices $x,y\in X_{0}$ with $x^{iM},y^{iM}\in V(Q^{i}_{0})$. From (i) we know that $M^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. We argue that (at least) one of $M^{i}_{0}\cup\{xx^{iM}\}$ or $M^{i}_{0}\cup\{yy^{iM}\}$ contains no (covered near) half-layers of $Q^{i}_{0}$, which implies that we can define $u:=x$ or $u:=y$, respectively. Suppose for the sake of contradiction that this is false, i.e., both $M^{i}_{0}\cup\{xx^{iM}\}$ and $M^{i}_{0}\cup\{yy^{iM}\}$ contain a (covered near) half-layer of $Q^{i}_{0}$. As $M^{i}_{0}\cup\{xx^{iM}\}$ contains a (covered near) half-layer, we obtain that $M^{i}_{0}$ contains a 2-near half-layer $L_{x}$ of $Q^{i}_{0}$, where $x$ is one of the extension vertices and the remaining three extension vertices are covered by $M$; see Figure 5. Symmetrically, as $M^{i}_{0}\cup\{yy^{iM}\}$ contains a (covered near) half-layer, we obtain that $M^{i}_{0}$ contains a 2-near half-layer $L_{y}$ of $Q^{i}_{0}$, where $y$ is one of the extension vertices and the remaining three extension vertices are covered by $M$. As all extension vertices of $L_{x}$ and $L_{y}$ except $x$ or $y$, respectively, are covered by $M$, we obtain that $L_{x}\cap L_{y}=\emptyset$. On the other hand, by Lemma 15 (iii), $L_{x}$ and $L_{y}$ have the same direction. This implies that there is a direction $j\in[d]\setminus\{i\}$ such that $|M^{j}_{-}|\geq|L_{x}|+|L_{y}|=2\cdot(2^{d-3}-2)\geq 12$ and $|M^{i}_{-}|\leq 2\cdot 4=8$. In fact, as $z,x,y\notin V(M)$ the latter bound can be improved to $|M^{i}_{-}|\leq 5$. This however contradicts our rules (1)–(3) for choosing the direction $i$ (see property (iii)).
@@
4 Claim 4a: $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. This was argued in the two subcases before.
5 Claim 5a: If $N^{i}_{1}$ contains a half-layer of $Q^{i}_{1}$, then there is a vertex on $u^{i}$’s side of the half-layer that is avoided by $M$. From this claim it follows in particular that $N^{i}_{1}$ and $u^{i}$ satisfy property (H) in Theorem 8 for $Q^{i}_{1}$. To prove the claim, note that we have $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ (in case ai) or $N^{i}_{1}=M^{i}_{1}$ (in case aii), and therefore $N^{i}_{1}$ can only contain a half-layer of $Q^{i}_{1}$ if $M^{i}_{1}$ contains a half-layer of $Q^{i}_{1}$. Let $L$ be such a half-layer of $Q^{i}_{1}$ contained in $M^{i}_{1}$, and let $j\in[d]\setminus\{i\}$ be the direction of $L$. By (ii), this means that $M^{i}_{-}$ contains a (covered near) quad-layer of $Q_{d}$; see Figure 6. Lemma 15 (i) yields that $M^{i}_{1}$ contains no half-layers of $Q^{i}_{1}$ in a direction different from $j$, and thus $L$ is the only half-layer of $Q^{i}_{1}$ in $M^{i}_{1}$. If $u^{i}\in V(L)$, then $N^{i}_{1}=M^{i}_{1}\setminus\{u^{i}u^{iM}\}$ contains no half-layers, and there is nothing to show. If $u^{i}\notin V(L)$, then the vertex $u^{i}$ together with all vertices in $X_{1}$ lie in either $(Q^{i}_{1})^{j}_{0}$ or $(Q^{i}_{1})^{j}_{1}$. As $X_{1}$ is nonempty, this proves the claim.
@@
Case b: $X_{0}=\emptyset$. We define $Y:=\{x\in V(Q^{i}_{0})\mid x\text{ is odd and }xx^{i}\notin M\}$, i.e., these are odd vertices in $V(Q^{i}_{0})$ for which the incident edge in direction $i$ is not in $M$. Clearly, we have $z\notin Y$ as $z=\emptyset$ is even. Furthermore, as $X_{0}=\emptyset$ and $M$ is assumed to satisfy property (H), we know that $Y\neq\emptyset$.
@@
    If $M^{i}_{1}$ contains a near half-layer $L^{\prime}$ of $Q^{i}_{1}$, then we select $u\in Y$ as a vertex such that $u^{i}$ is an (even) end vertex of one of the edges of $L^{\prime}$;
@@
4 Claim 4b: $N^{i}_{0}$ contains no (covered near) half-layers of $Q^{i}_{0}$. Note that if $N^{i}_{0}=M^{i}_{0}$ or $N^{i}_{0}=M^{i}_{0}\setminus\{uu^{M}\}$, then the claim follows directly from (i). It remains to consider the case $N^{i}_{0}=(M^{i}_{0}\setminus\{uu^{M}\})\cup\{u^{M}u^{iM}\}$, which occurs if and only if $u^{M},u^{iM}\in V(Q^{i}_{0})$. In particular, rule (1’) does not apply, as it would imply $u^{iM}\in V(Q^{i}_{1})$. Using (i) again, it can be ruled out that $N^{i}_{0}$ contains a half-layer of $Q^{i}_{0}$, and it remains to consider covered near half-layers. If $N^{i}_{0}$ contains a covered near half-layer $L^{\prime}$, then by (i) $u^{M}u^{iM}$ is one of its edges. Therefore $M^{i}_{0}$ contains the 2-near half-layer $L^{\prime\prime}:=L^{\prime}\setminus\{u^{M}u^{iM}\}$; see Figure 7 (a). Let $j\in[d]\setminus\{i\}$ be the direction of $L^{\prime\prime}$, which is also the direction of the edge $u^{M}u^{iM}$. By Lemma 15 (iii), $M^{i}_{0}$ contains no 2-near half-layers of $Q^{i}_{0}$ in a direction different from $j$, so only $L^{\prime\prime}$ and possibly a second 2-near half-layer in direction $j$ qualify for application of rule (2’). However, the edge $uu^{M}$ removed by rule (2’) also has the direction $j$. This is a contradiction, however, as the edges $uu^{M}$ and $u^{M}u^{iM}$ have different directions (as $u\neq u^{iM}$), proving the claim.